Clopack Company manufactures one product that goes through one processing department called Mixing. All raw materials are introduced at the start of work in the Mixing Department. The company uses the weighted-average method of process costing. Its Work in Process T-account for the Mixing Department for June follows (all forthcoming questions pertain to June):

Work in Process—Mixing Department
June 1 balance   28,000   Completed and transferred to Finished Goods   ?
Materials   120,000        
Direct labor   79,500        
Overhead   97,000        
June 30 balance   ?        

The June 1 work in process inventory consisted of 5,000 units with $16,000 in materials cost and $12,000 in conversion cost. The June 1 work in process inventory was 100% complete with respect to materials and 50% complete with respect to conversion. During June, 37,500 units were started into production. The June 30 work in process inventory consisted of 8,000 units that were 100% complete with respect to materials and 40% complete with respect to conversion.

Required:
1. Prepare the journal entries to record the raw materials used in production and the direct labor cost incurred.

2. Prepare the journal entry to record the overhead cost applied to production.


3. How many units were completed and transferred to finished goods during the period?

4. Compute the equivalent units of production for materials.
`42.500
5. Compute the equivalent units of production for conversion.
6. What is the cost of beginning work in process inventory plus the cost added during the period for materials?
7. What is the cost of beginning work in process inventory plus the cost added during the period for conversion?
8. What is the cost per equivalent unit for materials?
9. What is the cost per equivalent unit for conversion?
10. What is the cost of ending work in process inventory for materials?
11. What is the cost of ending work in process inventory for conversion?
12. What is the cost of materials transferred to finished goods?
13. What is the amount of conversion cost transferred to finished goods?
14. Prepare the journal entry to record the transfer of costs from Work in Process to Finished Goods.
15-a. What is the total cost to be accounted for?

15-b. What is the total cost accounted for?
Answer #1

1) Journal entry

No General Journal Debit Credit
a Work in process-Mixing 120000
Raw material 120000
(To record raw material used)
b Work in process-Mixing 79500
Wages payable 79500
(To record direct labor)

2) Journal entry

No General Journal Debit Credit
a Work in process-Mixing 97000
Manufacturing overhead 97000
(To record overhead applied)

3) Unit completed and transferred out = 5000+37500-8000 = 34500

4) Equivalent unit of material = 42500

5) Equivalent unit of conversion = 34500+(8000*40%) = 37700

Note: Pls post whole que in the group of 4-4 sub parts as per HOMEWORKLIB RULES